OpenStudy (bella.c.nance):

Yanni through his paper airplane 15 1/2 feet. Adrian through his paper airplane 3/4 of Yanni's distance. What is the distance Adrian through his paper airplane? Explain how you found your answer.

jhonyy9 (jhonyy9):

so you need to calculi the 3/4 part of 15 1/2 feet to get the Adrian through his paper airplane 15 1/2 = (2*15+1)/2 = 31/2 so the 3/4 part of 31/2 will be x so x = 3/4 *31/2 = (3*31)/(4*2) = 93/8 x = 93/8 so Adrian through his paper airplane has a distance of 93/8 feet what is equal 93/8 = 11 5/8 hope helped